A human resources manager is investigating the effectiveness of a new employee orientation program. Three different instructors each train 8 new employees about the company and its policies, and then a test is given to all new personnel. The manager wants to know if differences in test score can be attributed to instructor at an alpha of .05. She observes the following results in SPSS:

Mean

Standard Deviation

Instructor A

76.75

8.067

Instructor B

84.13

8.871

Instructor C

90.75

6.964

Total

83.88

9.625


Sum of Squares

df

Mean Square

F

Sig.

Between Groups

784.750

2

392.375

6.122

.008

Within Groups

1345.875

21

64.089

Total

2130.625

23



The null hypothesis that all three group means are equal is rejected. She wants to do a Tukey HSD to determine how the means differ from one another, but her computer freezes. You offer to do it for her by hand.

1. The (closest) tabled value for the studentized range statistic (q) is ______ (to 2 decimals).

2. The appropriate value for the MSW (aka MSE) is ________ (to 3 decimals).

3. The appropriate value for n is ________ .

4. The HSD value is _______ (to 2 decimals).
Answer #1

1)

critical q with 0.05 level and k=3, N-k=28 df= 3.50

2)

from given table MSW =64.089

3)

. The appropriate value for n is =8

4)

Tukey's (HSD) =(q/√2)*(sp*√(1/ni+1/nj)         = 9.91